# OptiRepair: Closed-Loop Diagnosis and Repair of Supply Chain Optimization Models with LLM Agents

### v1

@@ 1 Introduction
OptiRepair splits optimization model diagnosis and repair into two independently deployable phases. Phase 1 iterates on solver IIS feedback to diagnose and repair constraint conflicts (up to 20 turns); it applies to any LP, not just supply chains. Phase 2 applies a domain oracle that checks whether the solution is operationally sound; we implement it for supply chains with five theory-grounded rationality checks (Section 3.3.2). Each phase is independently trainable, but combined they connect optimization (“is the model solvable?”) to operations (“does the solution make sense?”). Two evaluation tracks test the framework: OptiRepair-SC runs the full two-phase pipeline on 976 supply chain problems (10 error types, ME-1ME-10); OptiRepair-Debug isolates Phase 1 on $\sim$}5,000 generic LP problems (9 error types, four OR domains). This work extends a solver-in-the-loop evaluation (Ao et al., 2026d) with supply chain specialization, two-phase rationality validation, and a training pipeline benchmarked against 22 frontier API models (Table 2).
@@
Table 1 summarizes the central result. Section 2 reviews related work; Section 3 defines the evaluation framework; Section 4 details the training pipeline; Section 5 reports results; and Section 6 discusses implications and limitations.
@@ 2.2 LLM-Assisted Optimization
Multiple benchmarks now evaluate LLMs on optimization tasks. Most focus on one-shot formulation: NL4Opt (Ramamonjison et al., 2023), OptiBench (Yang et al., 2024), ORLM (Huang et al., 2025), MAMO (Huang et al., 2024), and others (Xiao et al., 2024; AhmadiTeshnizi et al., 2024). Several testbeds add multi-step interaction: PILOT-Bench (Ao et al., 2026c) evaluates LLM robustness under probabilistic tool failures and variable instruction quality, DP-Bench (Zhou et al., 2025) applies solver verification to dynamic programming formulation, and LEAN-LLM-OPT (Liang et al., 2026) coordinates multiple agents to auto-formulate large-scale problems via structured workflow; all target formulation accuracy rather than repairing faulty models. Tsui (2025) measured a 64.5% self-correction blind spot rate in LLMs; CorrectBench (Tie et al., 2025) benchmarks correction across reasoning domains but excludes operations research. SWE-bench (Jimenez et al., 2024) benchmarks iterative code debugging, but unit tests only sample program behavior, whereas a solver delivers complete, deterministic feedback. OptiRepair addresses this gap: it evaluates iterative diagnosis and repair of faulty models with solver feedback, extending the correctness criterion beyond feasibility to domain-specific operational rationality.
Table 2: Evaluation framework comparison. OptiRepair-SC is the first to evaluate both iterative diagnosis-and-repair and operational rationality in a closed-loop supply chain setting.
@@ 3.1 Multi-Echelon LP Formulation
subject also to demand propagation ($d_{n,t}=x_{n-1,t}$ for $n\geq 2$; $d_{1,t}=d_{t}$) and non-negativity constraints ($x_{n,t},\,I_{n,t},\,B_{n,t}\geq 0$); see Appendix A for the full specification. Constraint Eq. (2) enforces flow conservation at each echelon, while demand propagation transmits order variability upstream—the bullwhip effect (Lee et al., 1997; Chen et al., 2000; Simchi-Levi et al., 2008). Appendix A gives the complete formulation with all constraint families.
@@ 3.4 Evaluation Metrics
Among problems where the agent achieves Optimal, the fraction that also pass rationality: $\text{P2Pass}=\text{RRR}/\text{RR}$. A model with high RR but low P2Pass restores feasibility yet yields operationally nonsensical solutions. Decomposing $\text{RRR}=\text{RR}\times\text{P2Pass}$ separates repair skill from operational understanding; Section 5 reports all three metrics.
@@ 3.5 Problem Set Construction
OptiRepair-SC (OptiRepair Supply Chain) contains 976 instances across all ten error types (ME-1ME-10), with disjoint supply chain configurations between the training and test splits (Appendix G details the split). Each problem bundles a natural-language description, Gurobi Python code for the sabotaged LP, the ground-truth fix, and the error type label. Instances span 2–5 echelons, 12–24 periods, three demand patterns, holding-to-backorder ratios from 1:21{:}2 to 1:101{:}10, and both capacitated and uncapacitated variants. We validate every problem: the sabotaged code returns Infeasible (or Optimal-but-irrational for ME-5), and the ground-truth fix passes all applicable rationality checks.
OptiRepair-Debug.
To test whether Phase I transfers beyond supply chains, we include OptiRepair-Debug, which contains approximately 5,000 problems across nine generic LP error types (Types A–I) drawn from facility location, transportation, scheduling, and network flow models. OptiRepair-Debug exercises Phase I only—no Phase II oracle—confirming that the repair loop generalizes across domains without supply-chain-specific modifications. Section 5.2 summarizes OptiRepair-Debug results.
@@ 4 Training Pipeline
Two specialized 8B-parameter models drive the OptiRepair closed-loop pipeline—one per phase. The combined training procedure, OptiSTaR (Optimization Self-Taught Reasoning), cycles through STaR exploration seeded by solver feedback and GRPO refinement at each iteration. Section 4.1 motivates the two-model split; Section 4.2 details the iterative STaR procedure for Phase I repair; Section 4.3 covers oracle-based Phase II training; and Section 4.4 tests architecture generalization.
@@ 4.1 Two-Model Architecture
The OptiRepair pipeline uses two independently trained 8B-parameter models: a P1 model for Phase I repair (Infeasible $\to$ solver-feasible) and a P2 model for Phase II rationality repair (solver-feasible $\to$ operationally rational). Each model solves a self-contained problem with its own reward signal: P1 maximizes solver feasibility recovery using IIS feedback, while P2 maximizes oracle check compliance using domain-specific rationality criteria. Separating the two objectives isolates feasibility restoration from operational validation. Phase I demands LP structure and IIS-based diagnosis—skills that transfer across optimization domains (OptiRepair-Debug confirms this on four problem families). Phase II demands domain-specific knowledge (base-stock structure, bullwhip bounds, cost consistency) and can be swapped or extended without retraining Phase I. A joint model mixes both objectives: it can restore feasibility by deleting cost constraints, producing irrational solutions.
@@ 4.2 Iterative STaR for Phase I Repair
Phase I training uses an iterative self-taught reasoning (STaR) loop (Zelikman et al., 2022), cycling through beam search exploration (Wei et al., 2022), distillation, and reinforcement learning in three stages (Figure 5).
@@
Iterative Loop.
The STaR loop ran three iterations, seeded with 696 successful trajectories from GPT-5.2, o4-mini, and DeepSeek-R1 teachers. Beam trajectories for hard error types (ME-2, ME-4) are filtered against ground-truth rationality: trajectories whose repairs fail rationality checks are replaced with upsampled ground-truth traces.
Training Progression.
Three STaR iterations raise Phase I $\text{RR@}{5}$ from 44.4% to 75.7% and P1 RR to 97.2%.
The Indirect Role of Reinforcement Learning.
GRPO’s direct effect on greedy performance is marginal (+0.7pp in $\text{RR@}{5}$). Its value is indirect: GRPO refinement raises beam search solve rate from 81.4% to 98.3%, feeding higher-quality distillation data into the next STaR iteration. In multi-turn agentic settings, RL amplifies data quality rather than directly improving the greedy policy.
@@ 4.3 Oracle-Based Phase II Training
Phase II training teaches the model to recognize and repair operationally irrational solutions that remain mathematically feasible. The critical design choice is the rationality oracle—the automated verifier that defines “correct” Phase II behavior.
Oracle Design Determines Effectiveness.
Three oracle design decisions—error-type-aware check activation, artifact-check relaxation, and prefix-matched constraint operations—together determine Phase II effectiveness. Precise correctness rules are a prerequisite for effective training.
Solver-Based GRPO Reward.
Phase II GRPO grounds its reward in the solver: each rollout trajectory runs in Gurobi and the RationalityOracle scores it, tying the training signal to actual operational quality rather than textual similarity. The solver-based reward uses LoRA $\alpha=0.5$ because the signal is binary: the solution either passes rationality checks or fails. Paired with GRPO-trained Phase 1, the final pipeline reaches $\text{RRR}=81.7\%$ on the held-out test set.
Error-Type Specialization.
ME-5 errors (pure Phase II problems) reach 100% RRR. The hardest errors, ME-2 and ME-4, require both phases and yield the largest GRPO gains: from 6% and 27% RRR to 64.5% and 61.3%, respectively (Section 5.4).
@@ 4.4 Architecture Generalization
5 Experiments
Evaluation spans two tracks: OptiRepair-Debug (general OR repair, single-phase) and OptiRepair-SC (supply chain closed-loop, two-phase).
@@ 5.1 Experimental Setup
Evaluation tracks. OptiRepair-Debug tests single-phase repair on 450 stratified samples per model across nine error types (Types A–I), drawn from linear and integer programs spanning production planning, transportation, scheduling, and facility location. OptiRepair-SC tests the full two-phase closed-loop pipeline on 284 test problems (from the 976-problem set) across ten supply chain error types (ME-1ME-10), with a disjoint training split for domain-specialized models.
Models. The evaluation covers 30 models. OptiRepair-Debug tests 4 locally trained Qwen3-8B variants (SFT, OptiSTaR, DAPO, Curriculum) and 22 API models from seven families: OpenAI (7), Google (3), Anthropic Claude (4), DeepSeek (2), Qwen (4), Meta (1), and Moonshot (1). OptiRepair-SC tests the same 22 API models plus two trained 8B-parameter pipelines (Qwen3-8B and Llama-3.1-8B).
@@
5.2 OptiRepair-Debug Results: General OR Repair
OptiRepair-Debug tests whether Phase I’s domain-agnostic repair loop transfers beyond supply chains, spanning approximately 5,000 problems across 9 generic LP error types and 26 models. The trained Qwen3-8B-OptiSTaR reaches 95.3% RR@5 and 62.4% DA, exceeding the best API model (o4-mini at 86.2% RR@5) by 9.1 percentage points. Trained models also complete repairs in 2.25 steps on average versus 3.78 for the best API models. Full per-model and per-error-type breakdowns are available from the authors.
@@ 5.3 OptiRepair-SC Results: Supply Chain Closed-Loop
The gap concentrates in Phase 1: API models average 27.6% RR versus 97.2% for trained models—a 69.6 percentage point gap (Figure 7). On OptiRepair-Debug, the best API models exceed 86% RR@5 on general OR problems; the drop to 28% on supply chain problems reflects harder repair challenges: longer constraint chains, multiple interacting echelons (Simchi-Levi et al., 2008), and domain-specific variable naming. Once feasibility is restored, the validation phase poses a lower barrier: the 19 format-compliant API models average 86.7% P2Pass, with several above 90%. The bottleneck is finding a feasible repair, not judging its operational quality.
@@ 6.1 Managerial Implications
Because Phase 1 (feasibility repair) is domain-agnostic—confirmed across four OR problem families—extending to a new domain requires only a new Phase 2 oracle encoding that domain’s correctness criteria. Organizations can share one repair model across domains while maintaining domain-specific validation.

### v2

@@ 1 Introduction
OptiRepair splits optimization model diagnosis and repair into two independently deployable phases. Phase 1 iterates on solver IIS feedback to diagnose and repair constraint conflicts (up to 20 turns); it applies to any LP, not just supply chains. Phase 2 applies a domain oracle that checks whether the solution is operationally sound; we implement it for supply chains with five theory-grounded rationality checks (Section 3.3.2). Each phase is independently trainable, but combined they connect optimization (“is the model solvable?”) to operations (“does the solution make sense?”). We evaluate the full pipeline on OptiRepair-SC, 976 supply chain problems spanning 10 error types (ME-1ME-10), and benchmark agentic performance against 22 frontier API models (Table 2).
@@
Table 1 summarizes the central result. Section 2 reviews related work; Section 3 defines the evaluation framework; Section 4 details the training pipeline; Section 5 reports agentic performance; and Section 6 discusses implications and limitations.
@@ 2.2 LLM-Assisted Optimization
Multiple benchmarks now evaluate LLMs on optimization tasks. Most focus on one-shot formulation: NL4Opt (Ramamonjison et al., 2023), OptiBench (Yang et al., 2024), ORLM (Huang et al., 2025), MAMO (Huang et al., 2024), and others (Xiao et al., 2024; AhmadiTeshnizi et al., 2024). Several testbeds add multi-step interaction: PILOT-Bench (Ao et al., 2026c) evaluates LLM robustness under probabilistic tool failures and variable instruction quality, DP-Bench (Zhou et al., 2025) applies solver verification to dynamic programming formulation, and LEAN-LLM-OPT (Liang et al., 2026) coordinates multiple agents to auto-formulate large-scale problems via structured workflow; all target formulation accuracy rather than repairing faulty models. Tsui (2025) measured a 64.5% self-correction blind spot rate in LLMs; CorrectBench (Tie et al., 2025) benchmarks correction across reasoning domains but excludes operations research. SWE-bench (Jimenez et al., 2024) benchmarks iterative code debugging, but unit tests only sample program behavior, whereas a solver delivers complete, deterministic feedback. OptiRepair addresses this gap: it evaluates iterative diagnosis and repair of faulty models with solver feedback, extending the correctness criterion beyond feasibility to domain-specific operational rationality. Ao et al. (2026d) introduce a solver-in-the-loop framework that benchmarks diagnostic accuracy of off-the-shelf LLMs on optimization debugging without closing the training loop. We measure end-to-end agentic performance—whether an agent can repair a formulation and validate the resulting solution—rather than diagnostic accuracy alone; apply the framework to multi-echelon supply chain models with theory-grounded rationality checks; and train specialized 8B-parameter agents through iterative STaR with solver-verified rewards.
Table 2: Positioning against optimization-agent benchmarks. OptiRepair-SC is the first to evaluate both iterative diagnosis-and-repair and operational rationality in a closed-loop supply chain setting.
@@ 3.1 Multi-Echelon LP Formulation
subject also to demand propagation ($d_{n,t}=x_{n-1,t}$ for $n\geq 2$; $d_{1,t}=d_{t}$) and non-negativity constraints ($x_{n,t},\,I_{n,t},\,B_{n,t}\geq 0$); Appendix A gives the complete formulation with all constraint families. Constraint (2) enforces flow conservation at each echelon, while demand propagation transmits order variability upstream—the bullwhip effect (Lee et al., 1997; Chen et al., 2000; Simchi-Levi et al., 2008).
@@ 3.4 Evaluation Metrics
Among problems where the agent achieves Optimal, the fraction that also passes rationality checks: $\text{P2Pass}=\text{RRR}/\text{RR}$. An agent with high RR but low P2Pass restores feasibility yet yields operationally nonsensical solutions. Decomposing $\text{RRR}=\text{RR}\times\text{P2Pass}$ separates repair skill from operational understanding; Section 5 reports all three metrics.
@@ 3.5 Problem Set Construction
OptiRepair-SC (OptiRepair Supply Chain) contains 976 instances across all ten error types (ME-1ME-10), with disjoint supply chain configurations between the training and test splits (Appendix G details the split). Each problem bundles a natural-language description, Gurobi Python code for the sabotaged LP, the ground-truth fix, and the error type label. Instances span 2–5 echelons, 12–24 periods, three demand patterns, holding-to-backorder ratios from 1:21{:}2 to 1:101{:}10, and both capacitated and uncapacitated variants. We validate every problem: the sabotaged code returns Infeasible (or Optimal-but-irrational for ME-5), and the ground-truth fix passes all applicable rationality checks. We report phase-wise performance on OptiRepair-SC through RR (Phase I recovery), P2Pass (Phase II conditional pass), and RRR (end-to-end success).
@@ 4 Training Pipeline
Two specialized 8B-parameter models drive the OptiRepair closed-loop pipeline—one per phase. The training procedure, OptiSTaR, cycles through solver-seeded STaR exploration and GRPO refinement at each iteration.
@@ 4.1 Two-Model Architecture
The OptiRepair pipeline uses two independently trained 8B-parameter models: a P1 model for Phase I repair (Infeasible $\to$ solver-feasible) and a P2 model for Phase II rationality repair (solver-feasible $\to$ operationally rational). Each model solves a self-contained problem with its own reward signal: P1 maximizes solver feasibility recovery using IIS feedback, while P2 maximizes oracle check compliance using domain-specific rationality criteria. Separating the two objectives isolates feasibility restoration from operational validation. Phase I demands LP structure and IIS-based diagnosis under multi-turn solver interaction. Phase II demands domain-specific knowledge (base-stock structure, bullwhip bounds, cost consistency) and can be swapped or extended without retraining Phase I. A joint model mixes both objectives: it can restore feasibility by deleting cost constraints, producing irrational solutions.
@@ 4.2 Iterative STaR for Phase I Repair
Phase I training follows an iterative STaR loop (Zelikman et al., 2022): beam search exploration (Wei et al., 2022), supervised distillation, and GRPO refinement (Figure 5).
@@
The STaR loop ran three iterations, seeded with 696 successful trajectories from GPT-5.2, o4-mini, and DeepSeek-R1 teachers. Beam trajectories for hard error types (ME-2, ME-4) are filtered against ground-truth rationality: trajectories whose repairs fail rationality checks are replaced with upsampled ground-truth traces. Over the full loop, Phase I $\text{RR@}{5}$ rises from 44.4% to 75.7% and P1 RR reaches 97.2%.
GRPO’s direct effect on greedy performance is marginal (+0.7pp in $\text{RR@}{5}$). Its value is indirect: refinement raises beam search solve rate from 81.4% to 98.3%, feeding higher-quality distillation data into the next STaR iteration. In multi-turn agentic settings, RL amplifies data quality rather than directly improving the greedy policy.
@@ 4.3 Oracle-Based Phase II Training
Phase II training teaches the model to repair operationally irrational solutions that remain mathematically feasible. Three oracle design decisions shape training effectiveness: error-type-aware check activation, artifact-check relaxation, and prefix-matched constraint operations.
Phase II GRPO grounds its reward in the solver: each rollout trajectory runs in Gurobi and the Rationality Oracle scores it, tying the training signal to operational quality rather than textual similarity. The reward is binary (pass/fail), using LoRA $\alpha=0.5$. Paired with GRPO-trained Phase I, the final pipeline reaches $\text{RRR}=81.7\%$ on the held-out test set. ME-5 errors (pure Phase II problems) reach 100% RRR; the hardest errors, ME-2 and ME-4, require both phases and yield the largest GRPO gains: from 6% and 27% to 64.5% and 61.3%, respectively (Section 5.3).
@@ 4.4 Architecture Generalization
5 Agentic Performance
Evaluation focuses on OptiRepair-SC, a supply chain closed-loop benchmark with two-phase interaction.
@@ 5.1 Experimental Setup
Benchmark. OptiRepair-SC tests the full two-phase closed-loop pipeline on 284 test problems (from the 976-problem set) across ten supply chain error types (ME-1ME-10), with a disjoint training split for domain-specialized models.
Models. The evaluation covers 24 models: 22 API models from seven families—OpenAI (7), Google (3), Anthropic Claude (4), DeepSeek (2), Qwen (4), Meta (1), and Moonshot (1)—plus two trained 8B-parameter pipelines (Qwen3-8B and Llama-3.1-8B).
@@ 5.2 OptiRepair-SC Results: Supply Chain Closed-Loop
The gap concentrates in Phase 1: API models average 27.6% RR versus 97.2% for trained models—a 69.6 percentage point gap (Figure 7). This low Phase 1 rate reflects harder repair challenges: longer constraint chains, multiple interacting echelons (Simchi-Levi et al., 2008), and domain-specific variable naming. Once feasibility is restored, the validation phase poses a lower barrier: the 19 format-compliant API models average 86.7% P2Pass, with several above 90%. The bottleneck is finding a feasible repair, not judging its operational quality.
@@ 6.1 Managerial Implications
Because Phase 1 (feasibility repair) is designed as a domain-agnostic solver-interaction module, extending to a new domain requires only a new Phase 2 oracle encoding that domain’s correctness criteria. Organizations can share one repair model across domains while maintaining domain-specific validation.